What is Pettijohn’s ROE? Given the following figures ($million):

3VdK2gWBIHAwCOC4dv2mAI4Kp3Xs7mDRCu1grlQU

Select one:

a. 8.54%

b. 8.99%

c. 9.44%

d. 9.91%

e. 10.41%
Answer #1

Total assets=Total liabilities+Total equity

Total equity=(42000-29400)=$12600

ROE=Net income/Total equity

=(1133/12600)

=8.99%(Approx).